# Into the Unknown: Self-Learning Large Language Models

### v3

@@ I Introduction
The self-learning LLM is particularly useful for having an up-to-date model without training a new one from scratch while also minimizing human involvement, greatly reducing development costs. It can be applied in application domains where new facts continuously come up. One example is personalized sentiment prediction and emotion recognition. When new users enter the system, the LLM would be able to perform self-learning only on data related to the new users. Another potential application is the knowledge exchange between LLMs, allowing the knowledge of one LLM to be learned by another LLM automatically.
@@ IX-A Models
TABLE I: Details of LLMs used in the experiments.
@@ IX-C Results
TABLE II: Experimental results.
@@ X-E Results
TABLE III: Model evaluation results before and after self-learning.
@@ Acknowledgment
This work was financed by (1) the National Science Centre, Poland, project no. 2021/41/B/ST6/04471; (2) the statutory funds of the Department of Artificial Intelligence, Wroclaw University of Science and Technology; (3) the Polish Ministry of Education and Science within the programme “International Projects Co-Funded”; (4) the European Union under the Horizon Europe, grant no. 101086321 (OMINO). However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor European Research Executive Agency can be held responsible for them.

### v4

@@ I Introduction
The self-learning LLM is particularly useful for having an up-to-date model without training a new one from scratch while also minimizing human involvement, greatly reducing development costs. It can be applied in application domains where new facts continuously come up. One example is personalized sentiment prediction and emotion recognition [3, 4, 5, 6, 7, 8, 9, 10]. When new users enter the system, the LLM would be able to perform self-learning only on data related to the new users. Another potential application is the knowledge exchange between LLMs, allowing the knowledge of one LLM to be learned by another LLM automatically.
@@ IX-A Models
TABLE I: Details of LLMs used in the experiments. ”Model Name” shows the model’s codename in our work, while ”HF Name” provides the model’s ID available on the HuggingFace platform.
@@ IX-C Results
TABLE II: Experimental results. Each row presents each model’s evaluation result using a particular method for identification of PiUs. ”Cur” indicates the Curiosity score, ”Kaw” indicates the Knowledge-Limit Awareness score, ”brev” indicates the brevity penalty, and ”SLC” is the Self-Learning Capability score.
@@ X-E Results
TABLE III: Mistral-Instruct evaluation results before and after self-learning. ”Baseline” shows the metric values before self-learning; ”LoRA” shows the metric values after self-learning using Plain LoRA; ”Dyn-Adapt” shows the metric values after self-learning using Dynamic-Adapter.
@@
TABLE IV: Phi-3-Mini evaluation results before and after self-learning. ”Baseline” shows the metric values before self-learning; ”LoRA” shows the metric values after self-learning using Plain LoRA; ”Dyn-Adapt” shows the metric values after self-learning using Dynamic-Adapter.
Dataset Metric Baseline LoRA Dyn-Adapt
$\mathbf{Q_{H}}$ Hallucination Score 0.73 0.38 0.42
ROUGE-Lsum 0.47 0.48 0.48
LLM-Judge 0.39 0.39 0.39
Wiki Perplexity 8.33 11.41 11.41
Alpaca ROUGE-Lsum 0.34 0.35 0.34
LLM-Judge 0.25 0.27 0.25
To ensure the generalizability of our findings, we performed one additional experiment using phi-3-mini as the base of the self-learning LLM, utilizing the corresponding output from the Oracle-Selected experiment. The results are presented in Table IV. Both Plain LoRA and Dyn-Adapt were able to greatly reduce the hallucination on $Q_{H}$ after self-learning, indicating more consistent answering of those questions. Despite the steep drop in hallucination, the changes in ROUGE-Lsum and LLM-Judge scores were very minuscule; this is because the initial model was already able to answer many questions at least partially correct before self-learning. Notably, the increase in perplexity on the Wiki dataset after self-learning was relatively higher compared to the previous experiment with mistral-instruct. This can be explained by the fact that phi-3-mini is a smaller model; small models tend to be more sensitive to minute changes in the parameters. Still, the perplexity increase is considered quite small. Manual inspection into the model’s responses did not reveal any degradation in the text generation quality.
@@ XI Possible Issues and Potential Extensions
Limitations of self-learning. The goal of self-learning is to enable more efficient model updates, especially in application domains where new information is constantly emerging. This approach excels in dynamic environments. However, it may offer little to no benefits in narrow, closed application domains, such as chatbots that are designed to answer questions about specific products or services. In such cases, the better approach is to simply prepare training data with complete facts of the domain from the beginning. Furthermore, self-learning heavily depends on the quality of the initial model. If the initial model is not sufficiently strong, it may generate nonsensical questions during self-learning. Because of this, it is advised to evaluate the model’s Self-Learning Capability before subjecting it to independent learning.
@@ Acknowledgment
This work was financed by (1) the National Science Centre, Poland, project no. 2021/41/B/ST6/04471; (2) the statutory funds of the Department of Artificial Intelligence, Wroclaw University of Science and Technology; (3) the Polish Ministry of Education and Science within the programme “International Projects Co-Funded”; (4) CLARIN ERIC – European Research Infrastructure Consortium: Common Language Resources and Technology Infrastructure (period: 2024-2026) funded by the Polish Minister of Science under the programme: ”Support for the participation of Polish scientific teams in international research infrastructure projects”, agreement number 2024/WK/01; (5) the European Union under the Horizon Europe, grant no. 101086321 (OMINO). However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor European Research Executive Agency can be held responsible for them.